# Bin Yang1,2, Alexandru Paul Condurache1,2

### v1

@@ 3.3 Post-Processing
After the augmented image being processed, 2D predictions from the network must be reprojected into 3D space using some post-processing technique. To align with the new inference framework with stacked predictions, we first propose an extension of the standard KNN method [24], termed  KNN Ensembling. In the post-processing phase, all sub-clouds are iteratively processed with KNN and votes are ensembled for every point from the full cloud to obtain final predictions. However, we found that the extension still faces the challenge in appropriately weighting contributions of nearest neighbors in 3D coordinates. In addition, the inference time is accumulated due to iterative process. To address the limitation, we propose a novel algorithm called Nearest Neighbors Range Interpolation (NNRI). Its pseudo-code is detailed in Algo. 1.
After applying softmax to the network output, we begin by kernelizing 2D predictions and range images using a pre-defined kernel size ($3\times 3$ in our experiments). Next, we assign each point’s nearest neighbors in 2D space with corresponding 2D coordinates and stack them along the sub-cloud dimension. The relative depth between each point and its neighbors is computed by taking the absolute difference in depth values. To extract valid data for interpolation, a threshold is needed to filter out distant neighbors. According to the prior knowledge [16, 20], using a constant threshold is sub-optimal due to differing point densities in LiDAR data: closer points are more likely to be affected by outliers due to high density, while farther points struggle to find valid neighbors due to sparsity. To fit this underlying geometry, the range value of each point is employed to determine its cut-off value. By normalizing the range using pre-computed mean and standard deviation, the cut-off value is derived from an exponential function, which approximates the relationship between point-sensor distance and density [21]. This approach simplifies computation by avoiding the costly nearest neighbor search in 3D space to calculate exact density values and adaptively assigns a threshold to each point. Once valid nearest neighbors are identified, they are normalized within the range of $[0,1]$ to compute interpolation weights. Finally, softmax scores of all 3D points are interpolated by the weighted sum of their nearest neighbors. NNRI is designed to effectively mitigate the ”many-to-one” issue inherently in range-view methods by leveraging distance-wise local neighborhood information in both 2D and 3D.
Define : $N=N_{max}$ sub-clouds.
@@
Output : Array $Labels$ with predicted labels for all points.
@@
2:    $S_{s}(n,h,w,k)\leftarrow\texttt{unfold}(I_{scores},k,pad)$
3:    $S_{r}(n,h,w,k)\leftarrow\texttt{unfold}(R_{ranges},k,pad)$
@@
12: $N_{valid}(n,p,k)\leftarrow\texttt{clamp}(N_{rel}(n,p,k),\texttt{max}=D(p)$)
13:    $W(n,p,k)=1-\texttt{Normalize}(N_{valid}(n,p,k))$
@@
15:    $Scores(p)=\sum_{i}^{k^{2}\times n}W(n,p,k)*N_{s}(n,p,k)$
16:    $Labels=\texttt{argmax}_{c\in C}(Scores(p))$
17:Return $Labels$
@@
3.4 Network Selection
In order to pursue enhancement in the segmentation accuracy while possibly maintaining the high efficiency of range-view-based approaches, we revisited prior works and selected three light-weight CNN-based networks (FIDNet [48], SalsaNext [7] and CENet [5]) for integration. To further test the effectiveness across different network architectures, we additionally deploy RangeViT [1], a network composed of a series of Vision Transformer blocks [9], in the experimental phase. Original RangeViT uses a trainable KPConv-based 3D projector to get the point-wise predictions. We replace it with our post-processing component to achieve the full integration of our framework and train the model from scratch.
@@ 4.1 Settings
Datasets We conduct experiments on two public LiDAR semantic segmentation datasets. SemanticKITTI [2] dataset [2] consists of 22 sequences captured with a 64-beam LiDAR sensor, encompassing 19 semantic classes. The dataset is split as follows: sequences 00 to 10 (excluding 08) are used for training, sequence 08 is reserved for validation, and sequences 11 to 21 are designated for testing. nuScenes dataset [3] comprises 1,000 driving scenes recorded in Boston and Singapore using a 32-beam LiDAR sensor, leading to a relatively sparse point cloud. After merging similar and infrequent classes, the dataset includes 16 distinct semantic classes. Implementation Details Prior works experimented mostly with the resolution of $64\times 2048$ for SemanticKITTI [5, 24, 7], and $32\times 960$ [18] or $32\times 2048$ [1] for nuScenes. In contrast, we reduce the azimuth resolution while increasing the projection rate in FLARES mode: resolutions of $64\times 512$ for SemanticKITTI and $32\times 480$ for nuScenes are fixed for the input and the full point cloud is split into up to 3 and 2 ($N_{max}$) sub-clouds during training and inference. We directly use custom configurations of prior works [48, 5, 7, 1] to train the networks in FLARES mode. For training the selected models (excluding RangeViT) on the nuScenes dataset, we standardize the hyperparameter set since no default configurations are provided. Specifically, we use the AdamW optimizer [22] along with a OneCycle scheduler [29], setting the maximum learning rate to $1e^{-3}$ and training for 150 epochs. All models are trained on four NVIDIA GeForce GTX 1080Ti in distributed mode. Evaluation Metrics Following prior works, we assess the performance using Intersection-over-Union (IoU) $\text{IoU}_{i}=\frac{TP_{i}}{TP_{i}+FP_{i}+FN_{i}}$ and Accuracy (Acc) $\text{Acc}_{i}=\frac{TP_{i}}{TP_{i}+FP_{i}}$ for each class $i$, and calculate the mean Intersection-over-Union (mIoU) and mean Accuracy (mAcc) across all classes. $TP_{i}$, $FP_{i}$, and $FN_{i}$ represent the true positives, false positives, and false negatives for class $i$, respectively.
@@ 4.2 Comparative Study
We compare FLARES with baseline models across two datasets. As shown in Tab. 1, all four networks see significant improvements: SalsaNext has mIoU gains of 5.3% on SemanticKITTI and 2.3% on nuScenes, FIDNet improves by 7.9% and 3.9%, and CENet by 3.3% and 3.1%. RangeViT, as a ViT-based network, also exhibits huge enhancement in performance, confirming FLARES’s generalizability across different architectures. This improvement is especially prominent for smaller, dynamic, and under-represented classes such as truck, motorcycle, bicycle, pedestrian and bicyclist. In Fig. 4, we present a visual comparison. Notably, with the support of FLARES, the network demonstrates improved accuracy in segmenting foreground objects. An exception arises with the motorcyclist class in SemanticKITTI, where IoU scores decrease compared to the baseline. Diving into the problem, this can be traced back to the extremely low occurrence of annotations for that class in the dataset. In standard training on low-resolution range images, this class already suffers from poor representation. In FLARES mode, the occurrence is further reduced by splitting of the point cloud. This accumulation of downsampling prevents the network from optimizing on that rare class effectively and lead to inferior performance. In contrast, the improvement on nuScenes is more consistent as class frequencies are better balanced. We regard this as a corner case when testing on an class-imbalanced dataset. As a future work to resolve the issue, we aim to explore 3D reconstruction techniques to generate real-world-like pseudo LiDAR point clouds for augmentation [4, 23].
In Tab. 2, we further compare FLARES-boosted networks with other state-of-the-arts from various modalities. Given that methods we select preserve relatively fewer number of parameters, FLARES helps to improve the segmentation accuracy, being comparable to other point- or voxel-based approaches that deploy much larger and deeper networks. In addition, they outperform others significantly in latency, achieving superior trade-offs in accuracy and efficiency.
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@ 4.3 Ablation Study
To perform the ablation study, we test with CENet [5] on val set of SemanticKITTI [2]. Training Schema As shown in Tab.3, we conduct a comprehensive evaluation to examine the trade-off between accuracy and efficiency under different training configurations. In the standard mode, a high-resolution range image is used as the input to reproduce baseline results. While training in the STR paradigm[18], there is a slight performance drop, the memory consumption is significantly reduced by partitioning the full point cloud through azimuth-wise grouping. The limited increase in latency is due to the fact that STR has to additionally unite the batched prediction along azimuth resolution and the post-processing is still performed on the high-resolution image. In contrast, FLARES achieves remarkable improvements, showing increases of 2.7% in mIoU, 2.1% in mAcc, and 45% acceleration in inference. Similar to STR, FLARES reduces memory consumption compared to the standard mode, further optimizing the efficiency of range-view semantic segmentation.
Method Input resolution mIoU mAcc Lat.
Standard $1\times 64\times 2048$ 64.8 77.2 44 ms
STR [18] $1(5)\times 64\times 480$ 64.3 76.8 41 ms
FLARES $1(3)\times 64\times 512$ 67.5 79.3 24 ms
Table 3: Performance with varying training schemes is evaluated on val set of SemanticKITTI [2]. Input resolution is formatted in $B\times H\times W$ (batch size, image width and height). For STR [18], we use the configuration yielding the best performance on CENet [5]: a full resolution of $64\times 1920$ split into four $64\times 480$ sub-images. For the fair comparison, all models are integrated with proposed components and trained from scratch.
Data Augmentation As presented in Fig. 5(a), various data augmentation methods can enhance the segmentation performance by large margins and our WPD+ performs the best among them, which increases baseline mIoU by 5.7% and mAcc by 2.9%. Mix3D [25] increases the contextual information per frame by fusing one scene to another, however, the model can still suffer from class imbalance. RangeAug [18] consists of 4 different range-wise operations to enrich the semantic and structural cues. From their experiments, it demonstrated that this augmentation technique is especially effective for attention-based networks [33, 9], which possess significantly higher model capacities and are more reliant on data diversity for optimal performance. Similar to their experimental results, applying this method to lightweight CNN-based networks has shown limited success in improvement.
(a) WPD+
(b) #Frames sampled in WPD+
(c) Inclusion of synthetic dataset
(d) MCF
Figure 5: a) Initialization of training with standard geometric data augmentation (GDA) and benchmark several state-of-the-art 3D augmentation techniques, including Mix3D [25], RangeAug [18], and original WPD [13]. b) Different number of sampled frames for best performance c) A comparative plot showing the IoU scores of top-rare classes in scenarios both with and without the inclusion of the synthetic dataset. As reference, the class frequencies in the val set are provided. d) The models are trained using two different input configurations: either a single range image derived from the full cloud (FC) or a range image generated from a sub-cloud (SC), and inferred in FLARES mode. Note that for all trained models in a) - d), we leverage KNN Ensembling during post-processing phase.
@@
WPD+ includes two tunable parameters: the number of sampled frames from the original dataset and the use of the synthetic dataset. To find out the best parameter set, we conduct two additional experiments. Fig. 5(b) shows that sampling 6 frames results in the optimal performance, while increasing the number of frames beyond this point leads to performance degradation. This is likely due to the model’s limited scalability, similar to the effects observed in case of RangeAug[18], in addition, pasting too many pixels belonging to specific classes will again corrupt the semantic balance. Furthermore, Fig. 5(c) illustrates that the synthetic dataset plays a key-role in refining semantic prediction of top-rare classes. Noteworthily, using the synthetic dataset is efficient and practical because this allows us to customize sensor configurations to align with the target dataset and to define specific objects within the scene for downstream applications without any labor cost.
In the next stage, we study how unoccupied pixels in the range image can affect the performance. Firstly, we use the full point cloud for training and sub-clouds for inference, as the decreasing azimuth resolution can already resolve the problem of low 2D occupancy (the first column in Fig. 5(d)). However, empirical results reveal a drop in 2D accuracy when inferring from sub-clouds. This is possibly because of the domain shift caused by the occupancy difference in range images. In the next trial, we use the sub-cloud as training input instead, but this directly limits the performance due to the occupancy reduction. To tackle this compounded challenge brought by FLARES, we introduce Multi-Cloud Fusion, an additional data augmentation step which fuses multiple sub-clouds through occupancy padding during projection phase. As shown in Fig. 5(d), using MCF exhibits the highest performance in the series, achieving 1% increase in IoU over the model trained on full cloud. As an alternative, we apply RangeIP [44], an interpolation-based augmentation technique, to enhance 2D occupancy in the range image, but it results in slight worse accuracy compared to the baseline.
(a) Ablation on Post-Processing
(b) Inference Time [ms]
(c) Ablation on Input Resolution
Figure 6: a) - b): Various post-processing techniques are applied to the same trained model. KNN [24] and NLA [48] are used on the single range image generated from the full point cloud, whereas KNN Ensembling operates on multiple range images derived from sub-clouds. For NNRI, we assess it using both full cloud and sub-clouds. c): Evaluation on different input resolutions and corresponding number of sub-clouds ($N_{max}$).
Post-Processing As introduced in Sec. 3.3, a simple LiDAR model is built to approximate the distance-density function of 3D points and compute cut-off values for valid neighbors extraction. To verify the necessity of the step, some qualitative results are provided in Fig. 7. As can be seen, the adaptive cut-off values refine semantic predictions by better accommodating objects with varying density scales.
@@
Next, we explore the impact of various post-processing techniques on segmentation performance in Fig. 6(a) and 6(b). Regarding conventional KNN [24] as the baseline, NLA [48] demonstrates similar performance in both accuracy and latency. In contrast, we deploy our approach (NNRI) in the standard mode as well and observe a significant improvement: inference time is cut nearly 16% compared to KNN, while mAcc and mIoU increase by 2.8% and 2%, respectively. Unlike KNN, NNRI avoids the computational cost of Gaussian kernel calculations for distance weighting and directly performs nearest neighbor searches on the range image instead of in 3D space, further reducing computational overhead. NNRI interpolates class-wise scores based on relative depths rather than directly voting on hard labels, relying more on weighted information from nearest neighbors, which is the major reason why it outperform other post-processing approaches.
Switching to FLARES mode, we first extend the standard KNN approach to KNN Ensembling, which iteratively gathers votes from all points in each sub-cloud and aggregates them for the final prediction. While this extension improves the accuracy, it comes at approximately doubled latency cost. Conversely, when NNRI is adapted, it consistently provides notable improvements in both efficacy and efficiency. As a reference, we included evaluation scores on 2D predictions (-- dashed lines in Fig. 6(a)), showing that FLARES with NNRI significantly narrows the accuracy gap between 2D and 3D predictions. This suggests that our approach effectively mitigates the ”many-to-one” problem, offering substantial gains in segmentation performance. Input Resolution To further explore the optimal results in FLARES mode, we test various input resolutions, as shown in Fig. 6(c). From the experimental results, we found that resolutions of $512$ and $640$ deliver the best mIoU scores. Increasing the azimuth resolution beyond this point causes a slight performance drop. Nevertheless, all tested configurations outperform the baseline (the first row in Tab.3), demonstrating the superb effectiveness of our approach on range-view LiDAR semantic segmentation.

### v2

@@ 3.3 Post-Processing
After images are processed by the network, their 2D predictions must be reprojected into 3D space using a post-processing technique. Conventional methods [34, 65] typically rely on hard voting from nearest neighbors to infer predictions for points. However, they are designed for the single range image and requires an iterative processing of all images when working on FLARES, which results in increased computational complexity. Furthermore, these approaches are limited in their ability to appropriately weight the contributions of each neighbor in 3D coordinates, thus leading to sub-optimal performance. To leverage the advantages of multi-range settings and overcome these limitations, we propose a novel algorithm called Nearest Neighbors Range Interpolation (NNRI). Its pseudo-code is shown in Algo. 1.
Define : NN sub-clouds.
@@
Output : Array LabelsLabels with predicted labels for all points.
@@
[…]
@@
12: $\mathbf{N}_{valid}(n,p,k)\leftarrow\texttt{clamp}(\mathbf{N}_{rel}(n,p,k),\texttt{max}=D(p)$)
[…]
@@
[…]
17:Return LabelsLabels
@@
After applying softmax to the network output, we begin by kernelizing 2D predictions and range images using a pre-defined kernel size ($3\times 3$ in our experiments). Next, we assign each point’s nearest neighbors in 2D space with corresponding 2D coordinates and stack them along the sub-cloud dimension. The relative depth between each point and its neighbors is computed by taking the absolute difference in depth values. To extract valid data for interpolation, a threshold is needed to filter out distant neighbors. According to the prior knowledge [22, 27], using a constant threshold is sub-optimal due to differing point densities in LiDAR data: closer points are more likely to be affected by outliers due to high density, while farther points struggle to find valid neighbors due to sparsity. To fit this underlying geometry, the range value of each point is employed to determine its cut-off value. By normalizing the range using pre-computed mean and standard deviation, the cut-off value is derived from an exponential function, which approximates the relationship between point-sensor distance and density [30]. This approach simplifies computation by avoiding the costly nearest neighbor search in 3D space to calculate exact density values and adaptively assigns a threshold to each point. Once valid nearest neighbors are identified, they are normalized within the range of $[0,1]$ to compute interpolation weights. Finally, softmax scores of all 3D points are interpolated by the weighted sum of their nearest neighbors.
Overall, NNRI is designed to effectively mitigate the ”many-to-one” issue inherent in range-view methods by leveraging distance-wise local neighborhood information in both 2D and 3D. Moreover, the new approach eliminates the need for sub-cloud concatenation, as used in STR [25], and processes all range images in parallel, which significantly reduces the inference time.
@@ 4.1 Settings
Datasets We conduct experiments on two public LiDAR semantic segmentation datasets. SemanticKITTI [3] dataset [3] consists of 22 sequences captured with a 64-beam LiDAR sensor, encompassing 19 semantic classes. The dataset is split as follows: sequences 00 to 10 (excluding 08) are used for training, sequence 08 is reserved for validation, and sequences 11 to 21 are designated for testing. nuScenes dataset [16] comprises 1,000 driving scenes recorded in Boston and Singapore using a 32-beam LiDAR sensor, leading to a relatively sparse point cloud. After merging similar and infrequent classes, the dataset includes 16 distinct semantic classes.
Networks We revisited prior works and selected three light-weight CNN-based networks: FIDNet [65], SalsaNext [10] and CENet [7]) for integration. To further test the effectiveness across heterogeneous architectures, we additionally deploy RangeViT [2], a network composed of a series of Vision Transformer blocks [13], in the experimental phase. Original RangeViT uses a trainable KPConv-based 3D projector to get the point-wise predictions. We replace it with our post-processing component to achieve the full integration of our framework and train the model from scratch.
Implementation Details Prior works experimented mostly with the resolution of $64\times 2048$ for SemanticKITTI [7, 34, 10], and $32\times 960$ [25] or $32\times 2048$ [2] for nuScenes. In contrast, we reduce the azimuth resolution while increasing the projection rate in FLARES mode: resolutions of $64\times 512$ for SemanticKITTI and $32\times 480$ for nuScenes are fixed for the input and the full point cloud is split into up 3 and 2 sub-clouds during training and inference, respectively. FLARES uses the same loss configurations of prior works [65, 7, 10, 2], including a cross-entorpy loss (focal loss for RangeViT) and Lovász–Softmax loss [4]. For training the selected models (excluding RangeViT) on the nuScenes dataset, we standardize the hyperparameter set since no default configurations are provided. Specifically, we use the AdamW optimizer [32] along with a OneCycle scheduler [43], setting the maximum learning rate to $1e^{-3}$ and training for 150 epochs. All models are trained on four NVIDIA GeForce GTX 1080Ti in distributed mode.
@@ 4.2 Comparative Study
We compare FLARES with baseline models across two datasets. As shown in Tab. 1, all four networks see significant improvements: SalsaNext has mIoU gains of 5.3% on SemanticKITTI and 2.3% on nuScenes, FIDNet improves by 7.9% and 3.9%, and CENet by 3.3% and 3.1%. RangeViT, as a ViT-based network, also exhibits huge enhancement in performance, confirming FLARES’s generalization across different architectures. This improvement is especially prominent for smaller, dynamic, and under-represented classes such as truck, motorcycle, bicycle, pedestrian and bicyclist. Notably, with the support of FLARES, the network demonstrates improved accuracy in segmenting foreground objects. For instance, as shown in Fig. 4, our method’s predictions exhibit significantly higher correctness compared to the baseline.
An exception arises with the motorcyclist class in SemanticKITTI, where IoU scores decrease compared to the baseline. Diving into the problem, this can be traced back to the extremely low occurrence of annotations for that class in the dataset. In standard training on low-resolution range images, this class already suffers from poor representation. In FLARES mode, the occurrence is further reduced by splitting of the point cloud. This accumulation of downsampling prevents the network from optimizing on that rare class effectively and lead to inferior performance. In contrast, the improvement on nuScenes is more consistent as class frequencies are better balanced. We regard this as a corner case when testing on an class-imbalanced dataset. As a future work to resolve the issue, we aim to explore 3D reconstruction techniques to generate real-world-like pseudo LiDAR point clouds for enhanced augmentation [5, 33].
In Tab. 2, we compare the performance of networks boosted by FLARES with other state-of-the-art approaches across various modalities. Despite using relatively fewer parameters, FLARES-enhanced models achieve segmentation accuracy comparable to other point- or voxel-based methods that deploy much larger and deeper networks. Moreover, our approach significantly outperforms these alternatives in terms of latency, achieving an excellent trade-off between accuracy and efficiency.
@@
ΥSalsaNextΥ     $63.3$     $94.7$ $52.9$ $55.7$ $57.3$ $50.2$ $65.5$ $70.9$ $13.0$ $92.6$ $\mathbf{69.0}$ $77.7$ $20.5$ $90.4$ $65.8$ $80.8$ $65.0$ $63.4$ $55.4$ $62.4$
ΥSalsaNextΥ     $\mathbf{64.8}$     $\mathbf{95.1}$ $\mathbf{55.5}$ $\mathbf{56.5}$ $\mathbf{60.1}$ $\mathbf{53.7}$ $\mathbf{69.6}$ $\mathbf{74.1}$ $11.4$ $\mathbf{93.0}$ $68.9$ $\mathbf{78.9}$ $20.4$ $\mathbf{91.1}$ $\mathbf{67.6}$ $\mathbf{82.0}$ $\mathbf{66.7}$ $65.0$ $\mathbf{58.1}$ $\mathbf{64.1}$
@@
ΥFIDNetΥ     $65.1$     $95.3$ $51.0$ $57.0$ $54.8$ $58.1$ $68.1$ $68.9$ $14.4$ $92.3$ $68.3$ $78.0$ $32.3$ $91.6$ $67.6$ $83.7$ $66.6$ $68.8$ $55.1$ $64.8$
ΥFIDNetΥ     $\mathbf{67.4}$     $\mathbf{95.8}$ $\mathbf{56.7}$ $\mathbf{60.7}$ $\mathbf{58.1}$ $\mathbf{60.3}$ $\mathbf{72.5}$ $\mathbf{72.9}$ $15.8$ $\mathbf{93.2}$ $\mathbf{69.2}$ $\mathbf{79.9}$ $\mathbf{34.2}$ $\mathbf{91.9}$ $\mathbf{69.0}$ $\mathbf{84.6}$ $\mathbf{68.7}$ $\mathbf{70.3}$ $\mathbf{59.9}$ $\mathbf{66.9}$
@@
ΥCENetΥ     $66.6$     $95.6$ $58.5$ $61.6$ $51.7$ $50.2$ $74.5$ $72.4$ $23.2$ $91.4$ $69.6$ $77.1$ $31.7$ $91.1$ $66.6$ $83.8$ $69.9$ $68.3$ $60.3$ $68.7$
ΥCENetΥ     $\mathbf{68.0}$     $\mathbf{95.9}$ $\mathbf{61.1}$ $\mathbf{62.1}$ $\mathbf{57.2}$ $\mathbf{59.0}$ $\mathbf{77.2}$ $\mathbf{74.2}$ $12.2$ $\mathbf{92.2}$ $\mathbf{69.9}$ $\mathbf{78.7}$ $\mathbf{32.9}$ $\mathbf{91.8}$ $\mathbf{68.8}$ $\mathbf{84.7}$ $\mathbf{71.3}$ $69.9$ $\mathbf{62.9}$ $\mathbf{70.3}$
@@
ΥRangeViTΥ     $\mathbf{66.1}$     $\mathbf{95.6}$ $\mathbf{56.3}$ $\mathbf{60.5}$ $\mathbf{52.4}$ $\mathbf{57.1}$ $\mathbf{72.0}$ $\mathbf{69.7}$ $16.0$ $91.6$ $\mathbf{71.1}$ 77.3{77.3} $\mathbf{32.7}$ $91.4$ $67.4$ $83.1$ $68.0$ $68.1$ $58.0$ $\mathbf{67.5}$
@@
ΥSalsaNextΥ     $\mathbf{74.5}$     $\mathbf{75.0}$ $\mathbf{34.6}$ $\mathbf{90.4}$ $\mathbf{90.0}$ $\mathbf{43.8}$ $\mathbf{79.4}$ $\mathbf{72.9}$ $58.8$ $\mathbf{65.8}$ $\mathbf{79.9}$ $\mathbf{96.5}$ $70.1$ $\mathbf{74.0}$ $\mathbf{73.9}$ $\mathbf{87.6}$ $\mathbf{85.6}$
@@
ΥFIDNetΥ     $\mathbf{76.6}$     $\mathbf{77.6}$ $\mathbf{43.5}$ $\mathbf{92.9}$ $\mathbf{88.1}$ $\mathbf{56.5}$ $\mathbf{79.5}$ $\mathbf{77.7}$ $\mathbf{65.3}$ $\mathbf{67.0}$ $\mathbf{83.1}$ $\mathbf{96.6}$ $\mathbf{72.8}$ $\mathbf{75.0}$ $\mathbf{74.5}$ $\mathbf{88.5}$ $\mathbf{86.8}$
@@
ΥCENetΥ     $\mathbf{76.8}$     $\mathbf{76.7}$ $\mathbf{45.2}$ $\mathbf{93.5}$ $\mathbf{90.3}$ $49.6$ $\mathbf{83.1}$ $\mathbf{78.1}$ $\mathbf{66.4}$ $\mathbf{69.0}$ $\mathbf{82.5}$ $\mathbf{96.6}$ $\mathbf{73.9}$ $\mathbf{75.1}$ $\mathbf{74.6}$ $\mathbf{88.3}$ $\mathbf{86.3}$
@@
ΥRangeViTΥ     $\mathbf{77.0}$     $\mathbf{76.7}$ $39.2$ $\mathbf{93.0}$ $\mathbf{92.0}$ $\mathbf{55.2}$ $\mathbf{81.6}$ $\mathbf{77.2}$ $64.9$ $\mathbf{70.9}$ $\mathbf{84.1}$ $\mathbf{96.8}$ $\mathbf{74.1}$ $\mathbf{75.6}$ $\mathbf{75.1}$ $88.6$ $86.7$
@@ 4.3 Ablation Study
To perform the ablation study, we test with CENet [7] on val set of SemanticKITTI [3] dataset.
STR [25]     FLARES     WPD+     MCF     NNRI     mIoU     Lat.
                    63.1     44 ms
✓                     62.6     41 ms
    ✓                 64.2     46 ms
    ✓     ✓             65.5     -
    ✓     ✓     ✓         65.9     -
    ✓     ✓     ✓     ✓     67.5     24 ms
Table 3: Full ablation study on SemanticKITTI [3] dataset. The first row denotes the baseline results trained in standard mode with high-resolution range images ($64\times 2048$). For STR [25], we split the full range image into 4 sub-images, each has the resolution of $64\times 512$. For all experiments without NNRI, we employ the conventional KNN [34] post-processing.
Component Design In Tab. 3, we evaluate the contributions of each component in FLARES. Starting with the baseline results, we observe that applying STR [25] slightly reduces mIoU with the limited improvement in latency. Incorporating FLARES into the framework then yields a significant boost in performance compared to the baseline. However, since KNN post-processing must be iteratively applied to multiple range images, the inference speed becomes lower. Next, our two data augmentation methods effectively addressing the exacerbated class imbalance and amplified noise introduced by FLARES. Consequently, this leads to a further increase of 1.7% in mIoU. Finally, replacing the standard post-processing with NNRI not only improves the overall performance but also achieves a remarkable speed-up in inference due to its ability to run parallel computations across multiple range images. Essentially, compared to the baseline performance, we achieve 4.4% mIoU gain and around 45% speed-up through the new framework.
Figure 5: a) Ablation on the number of sampled frames for WPD+. b) Comparative plot of IoU scores for top rare classes with and without the synthetic dataset, alongside the class frequency distribution in the validation set. For all inferences, KNN [34] is used as the post-processing approach.
@@
WPD+ To address the challenge of class imbalance within FLARES, WPD+ is incorporated as a data augmentation step. It includes two tunable parameters: the number of sampled frames from the original dataset and the inclusion of the synthetic dataset. To determine the optimal configuration, we conducted two ablation experiments. Fig. 5(a) shows that sampling 6 frames results in the optimal performance, while increasing the number of frames beyond this point leads to performance degradation. Fusing too many frames can introduce noise and redundant information that may overwhelm the model, for instance, excessive fusion might result in overlapping objects and distorted boundaries when too many new objects are pasted into the scene. Furthermore, Fig. 5(b) illustrates that the synthetic dataset plays a key-role in refining semantic prediction of top-rare classes. Notably, using the synthetic dataset is both efficient and practical, as it allows us to customize sensor configurations to align with the target dataset and define specific objects within the scene for downstream applications without incurring any labor cost.
Figure 6: Ablation on various post-processing approaches: KNN [34], NLA [65], and NNRI are first compared in a single-range setting. *NNRI denotes the multi-range variant of our post-processing method. For fair evaluation, we also extend KNN to operate on multiple range images (*KNN). Dotted lines indicate the metrics for 2D predictions.
Post-Processing We explore the performance of various post-processing techniques on both efficacy and efficiency in Fig. 6. Regarding conventional KNN [34] as the baseline, NLA [65] demonstrates similar performance in both accuracy and latency. In contrast, we deploy our approach (NNRI) in the standard setting as well and observe a significant improvement: inference time is cut nearly 16% compared to KNN, while mAcc and mIoU increase by 2.8% and 2%, respectively. Unlike KNN, NNRI avoids the computational cost of Gaussian kernel calculations for distance weighting and directly performs nearest neighbor searches on the range image instead of in 3D space, which further reduces computational overhead. NNRI interpolates class-wise scores based on relative depths rather than directly voting on hard labels, relying more on weighted information from nearest neighbors, which is the major reason why it outperform other post-processing approaches.
Next, switching to FLARES mode, we first implement an extension of KNN, which iteratively gathers votes from all points in each sub-cloud and aggregates them for the final prediction. While this extension improves the accuracy, it comes at approximately doubled latency cost. Conversely, when NNRI is adapted to all sub-clouds, it consistently provides notable improvements in both efficacy and efficiency. As a reference, we included evaluation scores on 2D predictions, showing that FLARES with NNRI significantly narrows the accuracy gap between 2D and 3D predictions. This suggests that our approach effectively mitigates the ”many-to-one” problem and offers substantial gains in segmentation performance.
@@
Furthermore, we optimize the implementation of NNRI by leveraging an adaptive cut-off value to filter valid nearest neighbors. This parameter is derived by approximating the internal LiDAR geometry and the distance-density relationship of 3D points. To illustrate its impact, qualitative results are provided in Fig. 7. As shown, the adaptive cut-off refines semantic predictions by better accommodating objects with varying density scales.
Figure 8: (a) Ablation on number of sub-clouds (NN) with fixed image resolution of $64\times 512$. (b) Ablation on image width $W$ (NN is fine-tuned for each configuration, respectively).
Input Configuration In Fig. 8(a), we investigate in the effect on the number of splitting. It shows that increasing the number of sub-clouds beyond 4 degrades performance due to insufficient occupancy and projection artifacts. Conversely, fewer sub-clouds speeds up the inference but compromises accuracy. Overall, our choice of 3 sub-clouds optimally balances efficiency and accuracy. In Fig 8(b), we further test various azimuth resolution of input images. The experimental results indicate that width of $512$ delivers the best mIoU scores. Increasing or decreasing the azimuth resolution beyond this point causes a slight performance drop.